# Rethinking the Sampling Criteria in Reinforcement Learning for LLM Reasoning: A Competence-Difficulty Alignment Perspective

### v2

@@ Abstract
Reinforcement learning exhibits potential in enhancing the reasoning abilities of large language models, yet it is hard to scale for the low sample efficiency during the rollout phase. Existing methods attempt to improve efficiency by scheduling problems based on problem difficulties. However, these approaches suffer from unstable and biased estimations of problem difficulty and fail to capture the alignment between model competence and problem difficulty in RL training, leading to suboptimal results. To tackle these limitations, this paper introduces Competence-Difficulty Alignment Sampling (CDAS), which enables accurate and stable estimation of problem difficulties by aggregating historical performance discrepancies of problems. Then the model competence is quantified to adaptively select problems whose difficulty is in alignment with the model’s current competence using a fixed-point system. Experimental results across a range of challenging mathematical benchmarks show that CDAS achieves great improvements in both accuracy and efficiency. CDAS attains the highest average accuracy against baselines and exhibits significant speed advantages compared to Dynamic Sampling, a competitive strategy in DAPO, which is 2.33 times slower than CDAS.
Figure 1: Average accuracy (left) and training GPU hours (right) of different sampling strategies.
1 Introduction
@@
Advanced large language models (LLMs) exemplified by DeepSeek-R1 guo2025deepseek and OpenAI O1 openaio1 demonstrate remarkable performance in challenging tasks like mathematics. As a core technology in their reports, the Reinforcement Learning algorithm (RL), such as Proximal Policy Optimization schulman2017ppo (PPO) and Group Relative Policy Optimization shao2024deepseekmathgrpo (GRPO), is employed to amplify the reasoning capabilities of the models. It works by utilizing a verifier as the reward model to guide the generation of high-quality reasoning chains without the need for data annotation. Despite the promise, the RL training is costly and hard to scale, particularly due to its low sample efficiency during the rollout phase. Recent studies yu2025dapo ; zeng2025simplerl ; bae2025online indicate that sampling overly difficult problems often results in no correct chains, while sampling overly simple problems contributes little to model capabilities, leading to computational waste. Consequently, a host of efforts are devoted to exploring sampling strategies for more efficient and stable RL training.
@@
Existing strategies draw inspiration from Curriculum Learning (CL) bengio2009curriculum ; narvekar2020curriculum , scheduling data based on problem difficulty to enhance training stability and efficiency. Curriculum Sampling Strategy team2025kimi relies on prior difficulty labels, which are excessively offline, neglecting the inherent capabilities of the model. Dynamic Sampling used by DAPOyu2025dapo demonstrates promising results by oversampling and filtering out problems with the pass rate equal to 1 and 0, which incurs substantial rollout overhead and compromises the training efficiency. Prioritized Sampling Strategy used by Kimi k1.5team2025kimi records the latest pass rate of each problem during training and adaptively assigns a higher sampling probability to those with lower pass rates. Overall, the pass rate has been widely adopted as a proxy for modeling problem difficulty.
@@
Figure 2: The variation in pass rate of problems in palin GRPO on Qwen-2.5 7Byang2024qwen2p5 using MATH dataset.
@@
However, these strategies tend to be suboptimal due to two main issues: (1) Unstable and Biased Estimations of Problem Difficulty Using Single-Step Pass Rate. As our experiment of training Qwen-2.5 7Byang2024qwen2p5 with MATH datasethendrycksmathdataset2021 shows (fig. 2), the pass rate of individual problems exhibits considerable fluctuations throughout the training process, leading to unstable estimations of problem difficulty, consistent with several prior workszheng2023secrets ; peng2023stabilizing . Moreover, focusing solely on the final pass rate can introduce a difficulty bias. For instance, the pass rate of problem 2 at step 5 happens to surpass that of problem 1, despite their distinct trajectories and difficulty levels. This demonstrates that the pass rate at a single step fails to capture the true complexity and learning dynamics associated with each problem. (2) Failing to Appropriately Capture the Alignment between Model Competence and Problem Difficulty. A common strategy, such as in curriculum sampling and prioritized sampling, is to assign higher sampling probabilities to more difficult problems with lower pass rates. A practical consequence is that overemphasizing difficult samples often leads to the selection of many zero-gradient problems with a pass rate of 0 in GRPO training, limiting the training efficiency. However, a more appropriate approach, as advocated in the curriculum learningbengio2009curriculum ; narvekar2020curriculum , is to prioritize problems that are more aligned with the current level of competency of the model for a more efficient and effective training.
@@
To address the issues above, we propose Competence-Difficulty Alignment Sampling (CDAS) for RL training, dynamically sampling problems whose difficulty matches the model competence at the step level. The core intuition behind CDAS is twofold: (1) instead of relying solely on the pass rate at a single step, an accumulative estimation that incorporates all historical information tends to yield a more stable assessment of problem difficulty; and (2) explicitly modeling model competence to measure its alignment with problem difficulty, thereby enabling more effective sampling decisions. Specifically, we model problem difficulty as the trajectory of performance discrepancy over training steps, where each point reflects the gap between the expected and actual pass rate. Then we use the centroid of this trajectory to provide a stable and accurate assessment of problem difficulty. Further, model competence is defined as the negative expected difficulty across all problems, and the absolute difference between model competence and a problem’s difficulty is used to quantify their alignment. Considering the dynamics of RL training, the competence-difficulty alignment estimation above is further formulated as a difficulty-based fixed-point system, which can iteratively converge and ensures the stability of training with theoretical guarantees.
To valid the effectiveness of CDAS, we conduct GRPO training on Qwen2.5-7Byang2024qwen2p5 using a series of sampling strategies team2025kimi ; yu2025dapo . Results across 7 comprehensive mathematical reasoning benchmarkshendrycksmathdataset2021 ; lewkowycz2022solvingminerva ; he2024olympiadbench ; cobbe2021traininggsm8k , show that CDAS consistently outperforms powerful baselines (see fig. 1), achieving the highest average accuracy of $46.77\%$. Compared to Dynamic Sampling, a highly competitive baseline, CDAS achieves slightly better performance $+0.12\%$ while considerably reducing the training step time overhead by $57.06\%$. Additionally, CDAS continually improves performance on more challenging benchmarks, achieving an accuracy of $11.77\%$ on AIME25, which surpasses the plain GRPO by $+4.79\%$ and Dynamic Sampling by $+2.19\%$. Subsequent analysis demonstrates that CDAS flexibly allocates computational resources by successfully sampling valuable problems that exhibit a pass rate not equal to 0 or 1. Further experiments on code generation tasks and larger 14B models also highlight the generalization ability of our method. The contributions of this paper can be summarized as follows:
@@
  • We identify and analyze the limitations of existing sampling strategies from a new perspective, highlighting the importance of stable difficulty estimation and dynamic competence-difficulty alignment in RL training for LLMs.
@@
  • We introduced Competence-Difficulty Alignment Sampling (CDAS), adaptively selecting problems that match the model competence, which is grounded in a theoretically guaranteed fixed-point system. Extensive experiments validate its effectiveness and efficiency.
2 Related Work
RL for LLMs reasoning. Reinforcement learning (RL) has been widely adopted to enhance the reasoning abilities of LLMs, especially in mathematics and programming tasks he2024olympiadbench ; trinh2024solving ; jain2025livecodebench ; penedo2025codeforces . Actor-critic-based methods, such as Proximal Policy Optimization (PPO) schulman2017ppo , utilize a value model to estimate the value function, guiding the policy updates. On the other hand, REINFORCE-based methods rely on policy gradients without a value model. Group Relative Policy Optimization (GRPO) shao2024deepseekmathgrpo normalizes rewards within a group of generated outputs, eliminating the need for a separate value model. REINFORCE++ hu2025reinforce++ enhances the classical REINFORCE algorithm williams1992simple by incorporating optimization techniques from PPO. Dynamic Sampling Policy Optimization (DAPO) yu2025dapo introduces several optimizations to enhance training efficiency and stability in long-CoT reasoning tasks.
Sampling Strategies for RL training. Effective sampling strategies are crucial for efficient RL training with LLMs. Coarse-grained curriculum learning team2025kimi ; xie2025logic gradually increases problem difficulty based on predefined labels. LIMR li2025limr introduces Learning Impact Measurement (LIM) to select problems that align with the model’s learning trajectory. Prioritized Sampling team2025kimi tracks the pass rate for each problem and samples problems in proportion to their failure rates. This approach directs the model’s focus toward more challenging problems. Dynamic Sampling yu2025dapo continues to sample problems within a batch until their pass rates are neither 0 nor 1. While this ensures that the model focuses on problems that are neither too easy nor too hard, it can lead to extremely heavy computational overhead as the training steps increase.
3 Preliminary
@@
Group Relative Policy Optimization GRPO utilizes group-based advantage without a value model, thereby reducing computational overhead. Formally, given a problem $x$, the correct answer $y$, and a group of sampled responses $\{\hat{y_{i}}\}_{i=1}^{G}$ with their corresponding rewards $\{r_{i}\}_{i=1}^{G}$, GRPO calculates the advantage by normalizing the rewards within each group. The original GRPO objective employs a sample-level loss calculation which potentially introduces length biases yu2025dapo ; liu2025understanding ; chu2025gpg , so we utilize a token-level policy gradient loss as our objective function:
$\mathcal{J}_{\textrm{GRPO}}(\theta)=\mathbb{E}_{[x\sim\mathcal{D},\{\hat{y_{i}}\}_{i=1}^{G}\sim\pi_{\textrm{old}}(\cdot|x)]}$ (1)
@@
$\frac{1}{\sum_{i=1}^{G}|\hat{y_{i}}|}\sum_{i=1}^{G}\sum_{t=1}^{|\hat{y_{i}}|}\bigg{(}\min\Big{(}r_{i,t}(\theta)\hat{A}_{i,t},\text{clip}\Big{(}r_{i,t}(\theta),1-\epsilon,1+\epsilon\Big{)}\hat{A}_{i,t}\Big{)}-\beta\mathbb{D}_{\text{KL}}[\pi_{\theta}\parallel\pi_{\text{ref}}]\bigg{)},$
@@
$r_{i,t}(\theta)=\frac{\pi_{\theta}(\hat{y}_{i,t}\mid x,\hat{y}_{i,<t})}{\pi_{\theta_{\text{old}}}(\hat{y}_{i,t}\mid x,\hat{y}_{i,<t})},\quad\hat{A}_{i,t}=\frac{r_{i}-\text{mean}(\{r_{i}\}_{i=1}^{G})}{\text{std}(\{r_{i}\}_{i=1}^{G})}.$ (2)
Rule-Based Reward The use of reward model usually leads to reward hacking problem gao2023scaling ; weng2024rewardhack , so we use a rule-based reward function. The reward is computed using the following rule:
@@
Notably, we do not employ a format reward in GRPO implementation. Prior research indicates that strict format constraints may limit the upper bound of model performance zeng2025simplerl ; singh2023beyond ; wang2024planning . Therefore, we directly use the final accuracy as the reward.
4 Competence-Difficulty Alignment Sampling
In this section, we will systematically elaborate on the framework of Competence-Difficulty Alignment Sampling (CDAS), including modeling the problem difficulty from the perspective of trajectory, measuring the alignment between model and problems, and how to incorporate our CDAS into iterative GRPO training as a fixed-point system.
4.1 Modeling Problem Difficulty as a Trajectory
@@
We start with a simple example in fig. 3, where the pass rate of a problem increases from 0 to 1 as the training step increases. Measuring difficulty based solely on the pass rate fails to provide a stable estimation due to the significant fluctuations in its pass rate along with the inherent limitations of information at a single step. Intuitively, by taking into account the historical performance across all previous samplings, we can obtain a much more stable and informative estimation. We plot the average historical pass rate up to the current step in fig. 3, which results in a much smoother curve that reflects the underlying historical trend.
Inspired by this, we turn to model the difficulty $D$ of a problem $x$ using the trajectory of the model’s performance discrepancy on the problem. Specifically, for the model $M_{n}$ at the $n$-th training step, its performance discrepancy $d$ on a $x$ is defined as
$d_{n}(x)=\hat{P}_{M_{n}}(y|x)-P_{M_{n}}(y|x),$ (4)
where $P_{M_{n}}(y|x)$ represents the actual probability of the model solving the problem $x$, estimated by the pass rate $s(x)$. $\hat{P}_{M_{n}}(y|x)$ represents the expected probability that the model can solve $x$. Then, given the existing trajectory $\{d_{1}(x),d_{2}(x),...,d_{n}(x)\}$, we quantify the difficulty $D$ of the problem using the centroid of the trajectory
$D_{n}(x)=\frac{1}{n}\sum_{k=1}^{n}\left(\hat{P}_{M_{k}}(y|x)-P_{M_{k}}(y|x)\right).$ (5)
The centroid considers the cumulative performance discrepancies over multiple training steps, providing a more accurate and robust measure of problem difficulty. The concept of performance discrepancy can be seen as a derivation from the $1-s(x)$ approach in Prioritized Samplingteam2025kimi , serving as a refined estimation of difficulty at each individual training step. The $1-s(x)$ metric can be regarded as a special case of the performance discrepancy, where the model is assumed to be capable of solving all problems. The introduction of $\hat{P}_{M_{n}}(y|x)$ amplifies the influence of steps where the actual pass rate deviates significantly from the model’s expectation.
The estimation of $\hat{P}_{M_{n}}(y|x)$ is supposed to account for both the problem difficulty and the model competence. Here we employ the sigmoid function, which is widely used in probability modeling, to describe such probability
$\hat{P}_{M_{n}}(y|x)=\frac{1}{1+e^{-(C_{n-1}-D_{n-1}(x))}}=\sigma(C_{n-1}-D_{n-1}(x)),$ (6)
where $C_{i-1}$ represents the model’s competence at the previous step, quantified as the negative expectation of the difficulties over all problems
$C_{n-1}=-\mathbb{E}_{x}[D_{n-1}(x)].$ (7)
When the model competence exceeds the problem difficulty, the expected probability of the problem being solved is higher. Otherwise, it is lower when the competence is lower than the difficulty.
4.2 Step-Level Adaptive Sampling based on Quantified Alignment
From eq. 6, we can see that $C_{n-1}-D_{n-1}(x)$ quantifies the degree of deviation between the model’s competence and the problem difficulty. As a result, we naturally introduce the definition of the alignment between a single problem $x$ and the model $M_{n}$ as
$\mathcal{A}(x,M_{n})=|C_{n-1}-D_{n-1}(x)|.$ (8)
The smaller the value of $\mathcal{A}$, the better the problem difficulty matches the current competence of the model. At each training step, we adopt a symmetric sampling method to construct the problem set $B=B^{-}\cup B^{+}$ from problems with the smallest $\mathcal{A}$ for rollout, where
$B^{-}$ $=\operatorname{argmin}_{|B|/2}\,\mathcal{A}(x,M_{n}),\quad\text{s.t.}\quad C_{n-1}-D_{n-1}(x)\leq 0,$ (9)
$B^{+}$ $=\operatorname{argmin}_{|B|/2}\,\mathcal{A}(x,M_{n}),\quad\text{s.t.}\quad C_{n-1}-D_{n-1}(x)>0.$ (10)
Compared to directly selecting the problems with the smallest $\mathcal{A}$, such symmetric method maintains a balanced proportion of hard and easy problems in $B$, promoting training stability.
4.3 Reinforcement Learning with CDAS

### v3

@@ Abstract
The low sampling efficiency during the rollout phase poses a significant challenge to scaling reinforcement learning for large language model reasoning. Existing methods attempt to improve efficiency by scheduling problems based on problem difficulties. However, these approaches suffer from unstable and biased estimations of problem difficulty and fail to capture the alignment between model competence and problem difficulty in RL training, leading to suboptimal performance. To address these challenges, we introduce Competence-Difficulty Alignment Sampling (CDAS). This approach allows for accurate and stable estimation of problem difficulties by aggregating historical performance discrepancies across problems. Subsequently, model competence is quantified to adaptively select problems whose difficulties align with the model’s current competence using a fixed-point system. Extensive experiments in mathematical RL training show that CDAS consistently outperforms strong baselines, achieving the highest average accuracy of 45.89%. Furthermore, CDAS reduces the training step time overhead by 57.06% compared to the widely-used Dynamic Sampling strategy, verifying the efficiency of CDAS. Additional experiments on different tasks, model architectures, and model sizes demonstrate the generalization capability of CDAS.
Codehttps://github.com/DeyangKong/CDAS
Introduction
@@
Advanced large language models (LLMs) exemplified by DeepSeek-R1 (Guo et al. 2025) and OpenAI O1 (Jaech et al. 2024) demonstrate remarkable performance in challenging tasks like mathematics. As a core technology in their reports, the Reinforcement Learning (RL) algorithm, such as Proximal Policy Optimization (Schulman et al. 2017) (PPO) and Group Relative Policy Optimization (Shao et al. 2024) (GRPO), is employed to amplify the reasoning capabilities of the models. It works by utilizing a verifier as the reward model to guide the generation of high-quality reasoning chains without the need for data annotation. Despite the promise, the RL training is costly and hard to scale, particularly due to its low sample efficiency during the rollout phase. Recent studies (Yu et al. 2025; Zeng et al. 2025; Bae et al. 2025) indicate that sampling overly difficult problems often results in no correct chains, while sampling overly simple problems contributes little to model capabilities, leading to computational waste. Consequently, a host of efforts are devoted to exploring sampling strategies for more efficient and stable RL training.
@@
Existing strategies draw inspiration from Curriculum Learning (CL) (Bengio et al. 2009; Narvekar et al. 2020), scheduling data based on problem difficulty to enhance training stability and efficiency. Curriculum Sampling Strategy (Team et al. 2025) relies on prior difficulty labels, which are excessively offline, neglecting the inherent capabilities of the model. Dynamic Sampling used by DAPO (Yu et al. 2025) demonstrates promising results by oversampling and filtering out problems with the pass rate equal to 1 and 0, which incurs substantial rollout overhead and compromises the training efficiency. Prioritized Sampling Strategy used by Kimi k1.5 (Team et al. 2025) records the latest pass rate of each problem during training and adaptively assigns a higher sampling probability to those with lower pass rates. Overall, the pass rate has been widely adopted as a proxy for modeling problem difficulty.
@@
Figure 1: Pass rate variations of two problems in plain GRPO on Qwen2.5-7B using MATH dataset.
@@
However, these strategies tend to be suboptimal due to two main issues: (1) Unstable and Biased Estimations of Problem Difficulty Using Single-Step Pass Rate. As our experiment of training Qwen2.5-7B (Yang et al. 2024) with MATH dataset (Hendrycks et al. 2021b) shows (Figure 1), the pass rate of individual problems exhibits considerable fluctuations throughout the training process, leading to unstable estimations of problem difficulty, consistent with several prior works (Zheng et al. 2023; Peng et al. 2023). Moreover, focusing solely on the final pass rate can introduce a difficulty bias. For instance, the pass rate of problem 2 at step 5 happens to surpass that of problem 1, despite their distinct trajectories and difficulty levels. This demonstrates that the pass rate at a single step fails to capture the true complexity and learning dynamics associated with each problem. (2) Failing to Appropriately Capture the Alignment between Model Competence and Problem Difficulty. A common strategy, such as in curriculum sampling and prioritized sampling, is to assign higher sampling probabilities to more difficult problems with lower pass rates. A practical consequence is that overemphasizing difficult samples often leads to the selection of many zero-gradient problems with a pass rate of 0 in GRPO training, limiting the training efficiency (Le et al. 2025). However, a more appropriate approach, as advocated in the curriculum learning (Bengio et al. 2009; Narvekar et al. 2020), is to prioritize problems that are more aligned with the current level of competency of the model for a more efficient and effective training.
@@
To address the issues above, we propose Competence-Difficulty Alignment Sampling (CDAS) for RL training, dynamically sampling problems whose difficulty matches the model competence at the step level. The core intuition behind CDAS is twofold: (1) instead of relying solely on the pass rate at a single step, an accumulative estimation that incorporates historical information tends to yield a more stable assessment of problem difficulty; and (2) explicitly modeling model competence to measure its alignment with problem difficulty, thereby enabling more effective sampling decisions. To realize these intuitions, we explicitly define two core concepts: Model Competence and Problem Difficulty. We then introduce Competence-Difficulty Alignment to quantify the gap between them, enabling us to sample problems that are optimally matched to the model’s current learning stage. To ensure this difficulty estimation is stable and robust against single-step fluctuations, we aggregate the problem’s historical performance over time. When integrated into the dynamics of RL training, this entire framework is formulated as a fixed-point system, which is theoretically guaranteed to converge, ensuring a stable and principled training process.
To validate the effectiveness of CDAS, we conduct mathematical reinforcement learning on Qwen2.5-7B (Yang et al. 2024) with GRPO, the most widely-used RL algorithm, and compare with extensive sampling strategies. The main findings are as follows: (1) Superior Performance and Efficiency: Results across six comprehensive mathematical reasoning benchmarks show that CDAS consistently outperforms powerful baselines, achieving the highest average accuracy of 45.89%. Compared to Dynamic Sampling, a highly competitive baseline, CDAS reduces the training step time overhead by 57.06% (see Figure 2(b)), verifying the efficiency of CDAS. (2) Efficient Sampling Mechanism: In-depth analysis indicates that CDAS successfully samples more problems that align with the model’s competence, reducing the proportion of zero-gradient problems, significantly less than other sampling strategies. (3) Proven Generalization and Scalability: Additional experiments on different tasks (i.e., code generation), model architectures (i.e., LLaMa), and model sizes (i.e., Qwen2.5-14B) demonstrate the generalization capability of CDAS.
The contributions of this paper can be summarized as follows:
@@
  • We identify and analyze the limitations of existing sampling strategies from a new perspective, highlighting the importance of stable difficulty estimation and dynamic competence-difficulty alignment in RL training.
@@
  • We introduce Competence-Difficulty Alignment Sampling (CDAS), adaptively selecting problems that match the model competence, which is grounded in a theoretically guaranteed fixed-point system. Extensive experiments validate its effectiveness and efficiency.
(a) Average Accuracy
(b) Total Training GPU Hours
Figure 2: Performance comparison of CDAS against baselines. CDAS achieves the best performance while demonstrating significant efficiency advantages compared to the strong Dynamic Sampling baseline.
Preliminary
@@
Group Relative Policy Optimization GRPO utilizes group-based advantage without a value model, thereby reducing computational overhead. Formally, given a problem $x$, the correct answer $y$, and a group of sampled responses $\{\hat{y_{i}}\}_{i=1}^{G}$ with their corresponding rewards $\{r_{i}\}_{i=1}^{G}$, GRPO calculates the advantage by normalizing the rewards within each group. The original GRPO objective employs a sample-level loss calculation which potentially introduces length biases (Yu et al. 2025; Liu et al. 2025; Chu et al. 2025), so we utilize a token-level policy gradient loss as our objective function:
@@
$\begin{split}&\mathcal{J}_{\textrm{GRPO}}(\theta)=\mathbb{E}_{[x\sim\mathcal{D},\{\hat{y_{i}}\}_{i=1}^{G}\sim\pi_{\textrm{old}}(\cdot|x)]}\\ &\quad\frac{1}{\sum_{i=1}^{G}|\hat{y_{i}}|}\sum_{i=1}^{G}\sum_{t=1}^{|\hat{y_{i}}|}\bigg(\min\Big(r_{i,t}(\theta)\hat{A}_{i,t},\\ &\quad\text{clip}\Big(r_{i,t}(\theta),1-\epsilon,1+\epsilon\Big)\hat{A}_{i,t}\Big)-\beta\mathbb{D}_{\text{KL}}[\pi_{\theta}\parallel\pi_{\text{ref}}]\bigg),\end{split}$ (1)
@@
$r_{i,t}(\theta)=\frac{\pi_{\theta}(\hat{y}_{i,t}\mid x,\hat{y}_{i,<t})}{\pi_{\theta_{\text{old}}}(\hat{y}_{i,t}\mid x,\hat{y}_{i,<t})},$ (2)
$\hat{A}_{i,t}=\frac{r_{i}-\text{mean}(\{r_{i}\}_{i=1}^{G})}{\text{std}(\{r_{i}\}_{i=1}^{G})}.$ (3)
Rule-Based Reward The use of reward model usually leads to reward hacking problem (Gao et al. 2023; Weng 2024), so we use a rule-based reward function. The reward is computed using the following rule:
@@
Notably, we do not employ a format reward. Prior research indicates that strict format constraints may limit the upper bound of model performance (Zeng et al. 2025; Singh et al. 2023; Wang et al. 2024). Therefore, we directly use the final accuracy as the reward.
Competence-Difficulty Alignment Sampling
In this section, we introduce the Competence-Difficulty Alignment Sampling (CDAS) framework. We first motivate the need for a dynamic and stable difficulty metric, then define model Competence, problem Difficulty, and their Alignment. Finally, we present the alignment-based sampling strategy and its convergence properties.
Rethinking Problem Difficulty: The Need for Relativity and Stability
The efficacy of advanced sampling strategies hinges on their ability to assess problem difficulty accurately. We argue that existing methods are limited by their failure to account for two aspects: the relativity and stability of difficulty.
The Relativity of Difficulty. Static curriculum learning strategies rely on a fixed, absolute measure of difficulty (e.g., predefined labels). However, problem difficulty is not an intrinsic constant; it is inherently relative to the competence of the model attempting to solve it. A problem that is challenging early in training may become trivial as the model improves, so a meaningful difficulty metric must dynamically reflect the evolving relationship between the model’s capability and the problem’s complexity.
@@
The Stability of Difficulty. Using the instantaneous pass rate as a difficulty proxy introduces high variance. As shown in Figure 3, pass rates fluctuate significantly despite an overall upward trend, making single-step observations noisy and unreliable. To achieve a more reliable estimation, it is necessary to aggregate the problem’s historical performance, smoothing out short-term noise to better capture the true learning trend.
These limitations often lead to misalignment between model competence and problem difficulty, resulting in overly hard or easy samples that yield zero-gradient updates. To address this, we propose a sampling method based on two core concepts: Model Competence and Problem Difficulty.
Quantifying the Core Concepts: Competence, Difficulty, and Alignment
To realize our vision of a dynamic and stable sampling method, we introduce three interconnected concepts to ensure a clear logical foundation.
Step 1: Defining Model Competence ($C$) We define the model’s competence $C_{n}$ as a measure of its overall ability to solve problems at the training step $n$. Formally, competence is the negative mean of problem difficulty across the dataset $\mathcal{D}$:
$C_{n}=-\mathbb{E}_{x\in\mathcal{D}}[D_{n}(x)]$ (5)
where $D_{n}(x)$ is the difficulty of problem $x$, which we will precisely define next. The negative sign indicates that as the average problem difficulty decreases (i.e., the model solves them more effectively), the model’s competence score $C_{n}$ correspondingly increases.
Step 2: Defining Problem Difficulty ($D$) With the macro-level concept of competence established, we can now define the difficulty of an individual problem. We distinguish between a problem’s volatile instantaneous difficulty ($d$) at a single step and its much more robust stable difficulty ($D$), which is informed by its history.
A problem’s instantaneous difficulty, $d_{n}(x)$, at step $n$ is defined as the gap between the model’s expected performance and its actual performance.
  • Actual Performance is measured by the model’s pass rate on problem $x$ at step $n$, denoted as $s_{n}(x)$.
  • Expected Performance, $\hat{P}_{M_{n}}(x)$, models the probability with which we anticipate the model will solve the problem. This expectation should naturally depend on the gap between the model’s prior competence, $C_{n-1}$, and the problem’s established difficulty, $D_{n-1}(x)$. A natural and widely used way to capture this dependency is through a Sigmoid transformation:
    $\hat{P}_{M_{n}}(x)=\sigma(C_{n-1}-D_{n-1}(x))$ (6)
    The Sigmoid function smoothly maps the real-valued competence–difficulty gap to a probability in $[0,1]$, a formulation also widely adopted in ability–difficulty modeling such as Item Response Theory and ELO rating systems (Baker 2001; Lord and Novick 2008). If competence far exceeds difficulty ($C\gg D$), the expected probability of success approaches 1. Conversely, if a problem is far too difficult for the model’s competence ($C\ll D$), the probability approaches 0. When competence and difficulty are balanced ($C\approx D$), the outcome is most uncertain, with a success probability of 0.5.
The instantaneous difficulty $d_{n}(x)$ is thus the difference between these two quantities:
$d_{n}(x)=\underbrace{\sigma(C_{n-1}-D_{n-1}(x))}_{\text{Expected Performance}}-\underbrace{s_{n}(x)}_{\text{Actual Performance}}$ (7)
If $d_{n}(x)>0$, the model underperformed relative to expectations, implying the problem was harder than anticipated at this step. If $d_{n}(x)<0$, the model overperformed, suggesting the problem was easier.
Finally, to achieve the desired stability, we define a problem’s stable difficulty, $D_{n}(x)$, as the centroid of its historical instantaneous difficulties, which smooths out noise over time:
$D_{n}(x)=\frac{1}{n}\sum_{k=1}^{n}{d_{k}}=\frac{n-1}{n}\cdot D_{n-1}(x)+\frac{1}{n}\cdot d_{n}(x)$ (8)
Step 3: Quantifying Competence-Difficulty Alignment ($\mathcal{A}$) With definitions for both competence ($C$) and difficulty ($D$), we can now naturally quantify their alignment for any given problem. We define the alignment, $\mathcal{A}$, as the absolute difference between the model’s competence and the problem’s difficulty:
$\mathcal{A}(x,M_{n})=|C_{n-1}-D_{n-1}(x)|$ (9)
This alignment value, $\mathcal{A}$, measures how closely a problem’s difficulty matches the model’s current learning frontier. A low $\mathcal{A}$ value indicates that a problem is well-matched to the model’s present capabilities, making it a prime candidate for efficient and effective training.
Alignment-based Sampling and Training
Symmetric Sampling Strategy A naive approach to leverage our alignment metric might be to simply sample problems with the lowest absolute value of $\mathcal{A}(x,M_{n})$. However, this could cause the sampling distribution to become overly biased. To foster a more balanced and robust training dynamic, we introduce a Symmetric Sampling Strategy. At each training step, we partition the entire problem set into two groups based on their relationship to the model’s current competence:
  • The Slightly Harder Group ($B^{+}$): Problems for which the difficulty is greater than the model’s competence ($D_{n-1}(x)>C_{n-1}$). These are problems that currently lie just beyond the model’s ability.
  • The Slightly Easier Group ($B^{-}$): Problems for which the difficulty is less than or equal to the model’s competence ($D_{n-1}(x)\leq C_{n-1}$). These are problems that are within the model’s current grasp.
We then construct the training batch $B$ by selecting the $|B|/2$ problems with the lowest alignment value $\mathcal{A}$ from each group, forming the final batch $B=B^{-}\cup B^{+}$. This balanced approach prevents the sampling process from being biased towards only one side. Further study of Symmetric Sampling Strategy is provided in Appendix E.
CDAS in Reinforcement Learning: A Stable Fixed-Point System We integrate CDAS framework into the GRPO training loop, as detailed in Algorithm 1. Since $|B|$ is usually much smaller than the size of the training set, performing a full update of problem difficulties at each step will lead to heavy computational overhead. Instead, for problem $x_{j}$, we record the number of times it is sampled as $t_{j}$ and update its difficulty only when it is sampled. Specifically, at the $i$-th step, the problem difficulty can be updated iteratively by
$D_{t_{j}}(x_{j})=\frac{t_{j}-1}{t_{j}}\cdot D_{t_{j}-1}(x_{j})+\frac{1}{t_{j}}\cdot d_{n}(x_{j})$ (10)
We set the initial difficulty $D_{0}(x)$ for all problems and the initial model competence $C_{0}$ to 0, signifying no prior system knowledge.